OpenStudy (anonymous):

The population of a town is 8000.If males are increased by 6% and females are increased by 10%, the population will be 8600. FInd the number of females in the town ?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

males x females 8000-x x*.06+(8000-x)*.1=8600-8000 800-600=.04x x=200/.04=5000 females=3000

OpenStudy (anonymous):

dipankarstudy- thnq, but please can u elobarate !! please.

OpenStudy (anonymous):

Male you consider it as X and Females are 8000-X Males increased to 6% means 6/100 and Females increased to 10% i.e. 0.1 is increased strength 600 By solving the equation you can get the value of X

OpenStudy (anonymous):

increase in the no. of males is x*6/100=.06x similarly increase in female is (8000-x)*10/100=.1(8000-x) by ur problem in total increase 8600-8000=600

OpenStudy (anonymous):

in ur first ans dipankarstudy, how did u get 800-600=.04x

OpenStudy (anonymous):

see (8000-x)*.1=800-.1x x*.06+(8000-x).1=x*.06+800-.1x=800-.04x

OpenStudy (anonymous):

awesome !! thnqq so much buddy !!